Hey there, I am looking for answers to these questions to review with my peers, thank you.
A butcher has some hamburger that is 19% fat and some that is 10% fat. How many pounds of each should be mixed to make 87 lb of hamburger that is 13% fat?
19% fat _________ ib
10% fat _________ ib

Answer by richwmiller(17219)

One problem at a time. No similar problems. You submitted three similar problems.

Don't waste your time nor my time with this nonsense (19% fat _________ ib
10% fat _________ ib) BTW What is an ib?
I'll set up the third one for you.
.19x+.10y=.13*87
x+y=87
solve for x
x=87-y
substitute
.19(87-y)+.10y=.13*87
Solve for y
Then
plug in
x=87-y and solve for x
The rest is left for you. Have fun.